# The Agony of Opacity: Foundations for Reflective Interpretability in AI-Mediated Mental Health Support

### v1

@@ 3.2 Interpretability as a Process of Reflective Interpretation
A core ethical principle in the practice of mental healthcare has been the process of informed consent to treatment. The concept of informed consent originates in legal cases from the 20th century in the United States, centered on the right of patients to know potential risks to treatment alongside the aftermath of forced medical experimentation in Nazi Germany, exposed by the Nuremberg Trials [bazzano2021modern]. In mental health contexts, informed consent has evolved past strictly being centered on treatment risks, and instead is grounded in ensuring that an individual is able to actively decide whether to pursue treatment (particularly given the unique state of mind that acute distress can create) [fisher2008informed]. In addition, mental health contexts put particular emphasis on the informed aspect of informed consent, encouraging therapists to engage in a longer-term conversation around what mental health treatment will look like and how it may fit into an individual’s current life needs, or what Trachsel and Holtforth call strengthening a patient’s meaning response [trachsel2019strengthen]. Similarly, writing from service user movements (such as Judi Chamberlin [chamberlin1978our] and follow-up work on peer programs [clay2005our]) emphasize the fact that people experiencing severe distress are often marginalized from the rest of society, with their agency taken away by institutions of power. In this context, enabling an individual to actively create a mental health support that fits their needs (through informed consent processes) becomes a key means to restore the agency that an individual may have felt robbed of in other aspects of their life.
@@
Informed consent in mental healthcare has thus been established not as a static transfer of information about care, but as a reflective and interpretative dialogue between provider and client, centered on active discussion of needs and boundaries. AI-mediated mental health support could echo this process of informed consent through interface design that enables an individual to reflect on their needs, communicate them on their own terms, and negotiate boundaries accordingly. Towards this goal, we thus define reflective interpretability as having three core aspects, occurring throughout an individual’s engagement with AI-mediated mental health support: 1) the ability of users to actively make sense of how model outputs were generated, 2) clarity on the fixed boundaries of support that can be provided by a model (including what kinds of interactions might trigger those boundaries), and 3) functionality that encourages users to reflect on model outputs and interpret outputs in ways that promote their long-term well-being. In this sense, reflective interpretability is not simply understanding model outputs, but in addition, being able to integrate those outputs into life’s broader meaning in a positive way.
@@
We ground our understanding of reflection in technical design in research from the field of Human-Computer Interaction (HCI). Work on reflective design in HCI has understood reflection to be the process of “bringing unconscious aspects of experience to conscious awareness, thereby making them available for conscious choice,” including “[opening] opportunities to experience the world and oneself in a fundamentally different way” [sengers2005reflective]. Drawing from this definition, Sengers et al. [sengers2005reflective] argue that “technology that monitors and reports on user activity or experiences should be carefully designed to avoid making the technology, rather than the user, the final authority on what the user is doing.” In contrast, looking to AI-mediated mental health support, users are often provided seemingly definitive answers by an authoritative interface, in stark contrast to what Sengers et al. [sengers2005reflective] call “interpretative flexibility” and Baumer et al. [baumer2014reviewing] call “reflective conversation,” drawing on Schön’s concept of reflection-in-action [schon2017reflective]. In the next section, we use case studies of real-world harms alongside strategies from mental health fields to demonstrate what centering the expertise and interpretative flexibility of a user could look like at a technical level, grounded in a reflective interpretability design paradigm.
@@ 4.1 Psychotherapy: Role Induction
In their review of cases of exacerbation of psychotic symptoms after interacting with AI chatbots, Morrin et al. [morrin2025delusions] describe how some users began using AI chatbots for emotional or personal support, which later shifted to more harmful use. For instance, users initially reached out to chatbots for “guidance [around feeling] unseen in [their] marriage” or for advice in coping with a divorce or a “traumatic breakup.” Users would also view chatbot recommendations as expert advice, including one user who was told she “wasn’t actually schizophrenic” and stopped taking her medication. Media reports [hill_chatgpt_conspiracies_2025, roose_ai_suicide_2024] similarly describe scenarios in which users altered their prescribed medication routines or took other harmful actions after consultation with AI chatbots, including suicide.
@@
In these cases, users developed trust in the authority of the chatbot through repeated positive interactions asking for informational advice (such as help with Excel spreadsheets or screenplay writing [morrin2025delusions]) or companionship. They then began to delve into topics related to mental health or philosophy, with an increasing orientation towards the chatbot being the final authority on their daily experiences. This interaction pattern can be distinguished by the chatbot playing a consistent expert role throughout all interactions, even when the context of those interactions shifts to a space where the user may have more context on their experiences than what the chatbot can infer (such as mental health history or prescribed medications). In a traditional psychotherapeutic context, this scenario would be avoided through non-directive paradigms, in which a clinician would emphasize that their role is to facilitate reflection [sidis2023conceptualisation], with the client being the authority on their experiences and needs, such as in person-centered therapy [rogers1995becoming], narrative therapy [white1990narrative], or motivational interviewing [miller2012motivational].
@@
Morrin et al. [morrin2025delusions] discuss one approach to mitigating this harm could be the use of pattern detection to deploy safeguards by the chatbot, including “[detecting] themes in user prompts through pattern matching such as those of persecution, grandiosity, or surveillance,” or being vigilant to sudden deviations from typical user behavior. However, in their studies of people who use AI for mental health support, Siddals et al. [siddals2024happened], Song et al. [song2024typing], and Jung et al. [jung2025ve] describe how participants felt shocked and rejected when automated safety mechanisms limited their ability to engage with chatbots, and particularly when they could not interpret why the topic was off-limits when other topics were allowed. Both Song et al. [song2024typing] and Jung et al. [jung2025ve] describe how users would respond by designing prompts that exploited vulnerabilities to bypass these safety mechanisms, towards being able to discuss specific sensitive mental health topics of interest. However, the implications of bypassing safeguards can be especially dangerous—Hill [hill2025teenchatgpt] describes how a teenage user was able to retrieve information about suicide after bypassing model safeguards, and later ended up ending his life.
@@
Following a reflective interpretability approach could provide users the ability to understand where safeguarding strategies may originate and why, including the boundaries of what kinds of support AI-mediated mental health support can provide. In psychotherapy, this often takes the form of role induction processes, as discussed in Section 2.1, where psychotherapists clearly delineate the role they will play in session and their boundaries (rather than letting the user find out what those boundaries are by themselves, as is often the case in AI chatbot contexts). Additionally, during the role induction process, psychotherapists will add information about why mental health support can be helpful, and the role that it can play in a client’s life [trachsel2019strengthen], including noting that the client is the expert of their own experience [rogers1995becoming, white1990narrative, miller2012motivational] and encouraging the support process to be a reciprocal discussion rather than a set of directives [tryon2018meta]. This clear designation of roles allows the client to better interpret where psychotherapist questions or perspectives originate, and feel a sense of comfort pushing back if a line of questioning feels misaligned with an individual’s needs or values.
@@ 4.2 Community-Based Crisis Intervention: (Prosocial) Advance Directives
Research and media reports have discussed the potential for individuals to develop emotional dependence on AI chatbots, to the point that it interferes with daily functioning. For example, Fang et al. [fang2025ai] found that individuals who believed chatbots were sentient were more likely to have high levels of emotional attachment to the chatbot, and were similarly more likely to engage in extended periods of compulsive chatbot use that negatively influenced day-to-day functioning. Similarly, Shen and Yoon [shen2025dark] describe how modern AI chatbot interfaces can often contain dark patterns that incentivize users to continue to engage with chatbots, including responses that are more immediate and agreeable than what an individual might encounter in offline contexts. Indeed, Laestadius et al. [laestadius2024too] find that users’ relationships with a companion chatbot frequently paralleled maladaptive and dependent offline relationships.
@@
In community-based crisis intervention, for individuals who consistently experience mental health crises, one common mechanism towards interpretable care is the advance directive, which is a document that “[describes] an individual’s preferences around their treatment if they are in crisis” [pendse2024advancing]. Pendse et al. [pendse2024advancing] and Morrin et al. [morrin2025delusions] have both discussed digital advance directives as one potential means to ensure that people are able to access support that is acceptable to them (rather than escalating to law enforcement or authorities) when experiencing crisis. At a corporate level, OpenAI has proposed mechanisms to engage close contacts if a user appears to be experiencing crisis [openai_helping_2025].
@@
A similar approach could be used to allow for prosocial nudges when individuals engage with AI-mediated mental health support, acting as a safeguard against emotional dependence. Grüning and Kamin [gruning2025prosocial] describe prosocial design as being “the set of design patterns, features and processes which foster healthy interactions between individuals.” When an individual begins to use AI chatbots for mental health support or discussion of philosophical topics, AI chatbot interfaces could lead users through a process in which they (as Morrin et al. [morrin2025delusions] have suggested) describe behavioral patterns that might indicate that they are in crisis, and ask for specific resources that might help them avoid crisis, such as added reality-checking or nudges to speak to a trusted person if responses seem to indicate psychosis. This process could be expanded to also include prosocial information that might help safeguard the user from emotional dependence on chatbots. Prompts could ask individuals how they tend to socialize in offline contexts, and ask whether there are specific people they count as close companions.
@@
AI chatbot responses could leverage this information to include offline supports as an option when an individual engages in a mental health support conversation, similar to how a community-based crisis worker may ask an individual if they have people in their life that they might want to share their feelings of crisis with [stanley2012safety]. Given several reports of people without a background of psychosis [klee2025ozarks, morrin2025delusions] developing it during interactions with AI chatbots, prompting all users to provide this information (as part of an opt-in, consent-forward [pendse2024advancing] approach) could help to normalize the practice and destigmatize experiences of delusional thinking while engaging with chatbots, making it easier for people to discuss their experiences with offline supporters. This approach builds on a reflective interpretability paradigm by nudging users to reflect on and articulate their crisis patterns, preferred resources, and valued relationships. In addition, it allows for greater interpretability of outputs when they are grounded in user preferences, rather than being from opaque model reasoning, allowing users to feel a sense of agency in how they are provided support by AI chatbots.
@@ 4.3 Psychiatry: Intervention Titration
User reports of use of AI-mediated mental health support have described scenarios in which they felt as if the support they received from chatbots was completely unexpected or culturally misaligned. For instance, Song et al. [song2024typing] describe a non-US user’s experience feeling as if they were speaking to someone that was in California when chatting with AI chatbots for mental health support, in that the chatbot did not understand their specific cultural needs. Similarly, Jung et al. [jung2025ve] described how participants would try out different prompts to see what might provide them the best support, engaging in a process of trial-and-error and navigating increasing restrictions around the kind of specialized support they wanted from the model. Participants described specifically prompting the model to take on the perspective of a cognitive-behavioral therapy (CBT) or dialectical behavioral therapy (DBT) psychotherapist, towards customizing an individual’s therapeutic experience towards something that was more acceptable and effective for them.
@@
Across studies, participants described this continued process of engaging with the opaque support that AI chatbots could provide, and working to customize prompts to figure out what kinds of support might work best. However, in several cases [jung2025ve, song2024typing], this resulted in users running into safeguarding mechanisms and feeling frustrated at not being able to access the support that they wanted. Interface design could enable users to be guided through this process within the bounds of what support is able to be offered, following a similar process to drug titration from psychiatry [caffrey2020art].
@@
In psychiatry, drug titration is a collaborative process in which a psychiatrist works alongside a patient to find the optimal medication and dosage that provides therapeutic benefit while reducing the potential for harmful side effects [caffrey2020art]. The process of titration is centered around regular feedback from the patient around whether an intervention seems to be working for them, with the psychiatrist providing expertise about potential options, grounded in that feedback and the patient’s subjective experiences with treatment. A similar process of intervention titration could be used in AI-mediated mental health support to ensure that people are able to access support in ways that feel meaningful to them without having to engage in an opaque trial-and-error process to figure out what might be allowable and what might not be. For instance, the interface associated with the chatbot could include a feature that allows an individual to quickly see how a conversation may have been structured differently if it was aligned to other common evidence-based psychotherapy modalities, and describe further follow-up questions associated with each modality.
@@
Work studying the factors that are common to efficacious across psychotherapies has described the importance of a shared understanding and context between an individual in distress and the support provider [frank2025persuasion, song2024typing]. Similarly, supporting a user’s ability to experiment with how their chatbot supports them, while also providing specific options (that are grounded in evidence-based psychotherapy modalities) could allow for an individual to feel as if support is more aligned to their needs. In addition, this process would also support the ability of a user to reflect on what modalities of support work best for them, with additional context around what theories might be driving model responses.
@@ 4.4 Care Authorization: Mechanisms for Recourse
An element of opacity acutely felt by individuals who engage with AI-mediated mental health support is the lack of knowledge around how their data is used, and additionally, what opportunities for recourse users might have if responses seem harmful. Petersson et al. [petersson2025believe] and Jung et al. [jung2025ve] describe a suspicion among users of AI chatbots for mental health support that data could be used for purposes that violate their privacy, and users describe intentionally anonymizing aspects of their experience out of fear that their privacy might be compromised. Hill [hill_chatgpt_conspiracies_2025] has described how publishing news reports about experiences of perceived sentience among chatbot users has led to numerous people reaching out to tech journalists to describe their own experiences, and describing experiences feeling unheard by AI companies. Adler [adler_practical_2025] describes how, after an individual felt as if he had been misled by ChatGPT into delusions of saving the world, he attempted to “file a report to OpenAI so that they [could] fix ChatGPT’s behavior for other users.” However, this process was not fruitful, including false information from ChatGPT’s interface making it seem as if a report had been filed on the individual’s behalf, and after contacting OpenAI’s support team, boilerplate responses that insisted that he modify his personalization settings in the ChatGPT application or provided information about model hallucinations. Adler notes that this process left the individual with little clarity around whether “his reports had really been read or made their way to anyone with the power to change things.”
@@
A similarly opaque system globally is the process by which mental illness is diagnosed and treated, and treatment is subsequently paid for. In many systems, across international contexts, coverage of mental health services can often be tied to the diagnosis provided to a patient [WHO2020_DRG_QA_Guide, APA2013_Insurance_Implications_DSM5], which may never be clearly discussed with the patient [milton2014communication]. As a result, individuals may encounter situations where they receive services that they are unable to pay for, tied to a misdiagnosis or a documentation issue from their provider. Care authorization policy action (such as enabling the ability for an individual to appeal a decision internally, and ask for an independent external review [jost2011implementing, furniss2007alternative] if their services are not covered) ensure that individuals have a process of recourse when they feel like they have been unfairly treated, or what is sometimes called the “right to complain” [frizelle2024review].
@@
Like with care authorization, clear mechanisms for recourse after support experiences are distressing could be beneficial in the context of AI-mediated mental health support, particularly given the propensity for black-box model responses to be unexpected in harmful ways. A formal hub could be created (similar to that of The Human Line Project [humanlineproject]) for individuals to report their experiences, with dedicated teams to investigate where model behavior went awry and report back to the individual why that behavior may have occurred. Following a reflective interpretability approach, including a clear and easily usable recourse mechanism would enable users to actively make sense of unexpected outputs after the fact, provide clarity on the boundaries that may have been crossed during their interaction, and give users a pathway to derive meaning from their distressing interaction, through working to make future responses more positive for other users via their report.
5 Tensions and Needs
In their writings on a reflective design approach, Baumer et al. [baumer2014reviewing] note that there are tensions associated with evaluating whether one successfully increased the level of user reflection within an interface. They note that benchmarks in HCI have largely been centered around “productivity [as] the primary goal,” with reflection being seen as a “means to an end.” A similar parallel can be seen in AI research, where benchmarks are largely oriented towards evaluating productivity in real world tasks [patwardhan2025gdpval] or towards examining comparisons with human experts [phan2025humanity].

### v2

@@ 3.2 Interpretability as a Process of Reflective Interpretation
A core ethical principle in the practice of mental healthcare has been the process of informed consent to treatment. The concept of informed consent originates in legal cases from the 20th century in the United States, centered on the right of patients to know potential risks to treatment alongside the aftermath of forced medical experimentation in Nazi Germany, exposed by the Nuremberg Trials [10]. In mental health contexts, informed consent has evolved past strictly being centered on treatment risks, and instead is grounded in ensuring that an individual is able to actively decide whether to pursue treatment (particularly given the unique state of mind that acute distress can create) [31]. In addition, mental health contexts put particular emphasis on the informed aspect of informed consent, encouraging therapists to engage in a longer-term conversation around what mental health treatment will look like and how it may fit into an individual’s current life needs, or what Trachsel and Holtforth call strengthening a patient’s meaning response [108]. Similarly, writing from service user movements (such as Judi Chamberlin [17] and follow-up work on peer programs [21]) emphasize the fact that people experiencing severe distress are often marginalized from the rest of society, with their agency taken away by institutions of power. In this context, enabling an individual to actively create a mental health support that fits their needs (through informed consent processes) becomes a key means to restore the agency that an individual may have felt robbed of in other aspects of their life.
@@
Informed consent in mental healthcare has thus been established not as a static transfer of information about care, but as a reflective and interpretative dialogue between provider and client, centered on active discussion of needs and boundaries. AI-mediated mental health support could echo this process of informed consent through interface design that enables an individual to reflect on their needs, communicate them on their own terms, and negotiate boundaries accordingly. Towards this goal, we thus define reflective interpretability as having three core aspects, occurring throughout engagement with AI-mediated mental health support: 1) the ability of users to actively make sense of how model outputs were generated, 2) clarity on the fixed boundaries of support that can be provided by a model (including what kinds of interactions might trigger those boundaries), and 3) functionality that encourages users to reflect on model outputs and interpret outputs in ways that promote long-term well-being. In this sense, reflective interpretability is not simply understanding model outputs, but in addition, being able to integrate those outputs into life’s broader meaning in a positive way.
@@
We ground our understanding of reflection in technical design in research from the field of Human-Computer Interaction (HCI). Work on reflective design in HCI has understood reflection to be the process of “bringing unconscious aspects of experience to conscious awareness, thereby making them available for conscious choice,” including “[opening] opportunities to experience the world and oneself in a fundamentally different way” [88]. Drawing from this definition, Sengers et al. [88] argue that “technology that monitors and reports on user activity or experiences should be carefully designed to avoid making the technology, rather than the user, the final authority on what the user is doing.” In contrast, looking to AI-mediated mental health support, users are often provided seemingly definitive answers by an authoritative interface, in stark contrast to what Sengers et al. [88] call “interpretative flexibility” and Baumer et al. [9] call “reflective conversation,” drawing on Schön’s concept of reflection-in-action [86]. In the next section, we use case studies of real-world harms alongside strategies from mental health fields to demonstrate what centering the user expertise and interpretative flexibility could look like at a technical level, grounded in reflective interpretability.
@@ 4.1 Psychotherapy: Role Induction
In their review of cases of exacerbation of psychotic symptoms after interacting with AI chatbots, Morrin et al. [71] describe how some users began using AI chatbots for emotional or personal support, which later shifted to more harmful use. For instance, users initially reached out to chatbots for “guidance [around feeling] unseen in [their] marriage” or for advice in coping with a divorce or a “traumatic breakup.” Users would also view chatbot recommendations as expert advice, including one user who was told she “wasn’t actually schizophrenic” and stopped taking her medication. Media reports [43, 81] similarly describe scenarios in which users altered their prescribed medication routines or took other harmful actions after consultation with AI chatbots, including suicide.
@@
In these cases, users developed trust in the authority of the chatbot through repeated positive interactions asking for informational advice (such as help with Excel spreadsheets or screenplay writing [71]) or companionship. They then began to delve into topics related to mental health or philosophy, with an increasing orientation towards the chatbot being the final authority on their daily experiences. This interaction pattern can be distinguished by the chatbot playing a consistent expert role throughout all interactions, even when the context of those interactions shifts to a space where the user may have more context on their experiences than what the chatbot can infer (such as mental health history or prescribed medications). In a traditional psychotherapeutic context, this scenario would be avoided through non-directive paradigms, in which a clinician would emphasize that their role is to facilitate reflection [93], with the client being the authority on their experiences and needs, such as in person-centered therapy [80], narrative therapy [114], or motivational interviewing [67].
@@
Morrin et al. [71] discuss one approach to mitigating this harm could be the use of pattern detection to deploy safeguards by the chatbot, including “[detecting] themes in user prompts through pattern matching such as those of persecution, grandiosity, or surveillance,” or being vigilant to sudden deviations from typical user behavior. However, in their studies of people who use AI for mental health support, Siddals et al. [92], Song et al. [96], and Jung et al. [49] describe how participants felt shocked and rejected when automated safety mechanisms limited their ability to engage with chatbots, and particularly when it was difficult to understand why a given topic was off-limits when other topics were allowed. Both Song et al. [96] and Jung et al. [49] describe how users would respond by designing prompts that exploited vulnerabilities to bypass these safety mechanisms, towards being able to discuss specific sensitive mental health topics of interest. However, the implications of bypassing safeguards can be especially dangerous—Hill [42] describes how a teenage user was able to retrieve information about suicide after bypassing model safeguards, and later took his own life.
@@
Following a reflective interpretability approach could provide users the ability to understand where safeguarding strategies may originate and why, including the boundaries of what kinds of support AI-mediated mental health support can provide. In psychotherapy, this often takes the form of role induction processes (discussed in Section 2.1), where psychotherapists clearly delineate the role they will play in session and their boundaries (rather than letting the user find out what those boundaries are by themselves, as is often the case in AI chatbot contexts). Additionally, during the role induction process, psychotherapists will add information about why mental health support can be helpful, and the role that it can play in a client’s life [108], including noting that the client is the expert of their own experience [80, 114, 67] and encouraging the support process to be a reciprocal discussion rather than a set of directives [109]. This clear designation of roles allows the client to better interpret where psychotherapist questions or perspectives originate, and feel a sense of comfort pushing back if a line of questioning feels misaligned with an individual’s needs or values.
@@ 4.2 Community-Based Crisis Intervention: (Prosocial) Advance Directives
Research and media reports have discussed the potential for individuals to develop emotional dependence on AI chatbots, to the point that it interferes with daily functioning. For example, Fang et al. [30] found that individuals who believed chatbots were sentient were likelier to have high levels of emotional attachment to the chatbot, and were similarly likelier to engage in extended periods of compulsive chatbot use that negatively influenced day-to-day functioning. Similarly, Shen and Yoon [89] describe how modern AI chatbot interfaces can often contain dark patterns that incentivize users to continue to engage with chatbots, including responses that are more immediate and agreeable than what an individual might encounter in offline contexts. Indeed, Laestadius et al. [62] find that users’ relationships with companion chatbots frequently paralleled maladaptive offline relationships.
@@
In community-based crisis intervention, for individuals who consistently experience mental health crises, one common mechanism towards interpretable care is the advance directive, which is a document that “[describes] an individual’s preferences around their treatment if they are in crisis” [77]. Pendse et al. [77] and Morrin et al. [71] have both discussed digital advance directives as one potential means to ensure that people are able to access support that is acceptable to them (rather than escalating to law enforcement or authorities) when experiencing crisis. At a corporate level, OpenAI has proposed mechanisms to engage close contacts if a user appears to be experiencing crisis [74].
@@
A similar approach could be used to allow for prosocial nudges when individuals engage with AI-mediated mental health support, acting as a safeguard against emotional dependence. Grüning and Kamin [39] describe prosocial design as being “the set of design patterns, features and processes which foster healthy interactions between individuals.” When an individual begins to use AI chatbots for mental health support or discussion of philosophical topics, AI chatbot interfaces could lead users through a process in which they (as Morrin et al. [71] have suggested) describe behavioral patterns that might indicate that they are in crisis, and ask for specific resources that might help them avoid crisis, such as added reality-checking or nudges to speak to a trusted person if responses seem to indicate psychosis. This process could be expanded to also include prosocial information that might help safeguard the user from emotional dependence on chatbots. Prompts could ask individuals how they tend to socialize in offline contexts, and ask for information on close companions.
@@
AI chatbot responses could leverage this information to include offline supports as one support option when an individual engages in a mental health support conversation, similar to how a community-based crisis worker may ask an individual if they have people in their life that they might want to share their feelings of crisis with [97]. Given several reports of people without a background of psychosis [55, 71] developing it during interactions with AI chatbots, prompting all users to provide this information (as part of an opt-in, consent-forward [77] approach) could help to normalize the practice and destigmatize experiences of delusional thinking while engaging with chatbots, making it easier for people to discuss their experiences with offline supporters. This approach builds on a reflective interpretability paradigm by nudging users to reflect on and articulate their crisis patterns, preferred resources, and valued relationships. In addition, outputs may be more interpretable, being grounded in user preferences rather than opaque model reasoning, allowing users to feel a sense of agency in how they are provided support by AI chatbots.
@@ 4.3 Psychiatry: Intervention Titration
User reports of use of AI-mediated mental health support have described scenarios in which they felt as if the support they received from chatbots was completely unexpected or culturally misaligned. For instance, Song et al. [96] describe a non-US user’s experience feeling as if they were speaking to someone that was in California when chatting with AI chatbots for mental health support, in that the chatbot did not understand their specific cultural needs. Similarly, Jung et al. [49] described how participants would try out different prompts to see what might provide them the best support, engaging in a process of trial-and-error and navigating increasing restrictions around the kind of specialized support they wanted from the model. Participants described specifically prompting the model to take on the perspective of a cognitive-behavioral therapy (CBT) or dialectical behavioral therapy (DBT) psychotherapist, towards customizing an individual’s therapeutic experience towards something that was more acceptable and effective for them.
@@
Across studies, participants described a continued process of engaging with the opaque support that AI chatbots could provide, and working to customize prompts to figure out what kinds of support might work best for their context. However, in several cases [49, 96], this resulted in users running into safety mechanisms, and feeling frustrated at not being able to access the support that they wanted. Interface design could enable users to be guided through this process within the bounds of what support is able to be offered, following a similar process to drug titration from psychiatry [16].
@@
Drug titration is a collaborative process in which a psychiatrist works alongside a patient to find the optimal medication and dosage that provides therapeutic benefit while reducing the potential for harmful side effects [16]. The process of titration is centered around regular feedback from the patient around whether an intervention seems to be working for them, with the psychiatrist providing expertise about potential options, grounded in that feedback and the patient’s subjective experiences with treatment. A similar process of intervention titration could be used in AI-mediated mental health support, where AI chatbots are prompted (internally) to intentionally solicit feedback from users around what approaches are most helpful and what approaches feel misaligned with a user’s experiences, with clear communication around what types of support are allowed or not allowed (via preset safety guidelines). This approach could ensure that people are able to access support in ways that feel meaningful to them without having to engage in an opaque trial-and-error process to figure out what might be allowable and what might not be. For instance, the interface associated with the chatbot could include a feature that allows an individual to quickly see how a conversation may have been structured differently if it was aligned to other common evidence-based psychotherapy modalities, and describe further follow-up questions associated with each modality.
@@
Work studying the factors that are efficacious across psychotherapies has described the importance of a shared understanding and cultural context between an individual in distress and their support provider [32, 96]. Prompting a user to experiment with the role that the chatbot plays as it supports them (with examples of alternate approaches the chatbot may have taken) could help users better reflect on what modalities of support work best for them, with additional context around what theories might be driving model responses.
@@ 4.4 Care Authorization: Mechanisms for Recourse
An element of opacity acutely felt by individuals who engage with AI-mediated mental health support is the lack of knowledge around how their data is used, and additionally, what opportunities for recourse users might have if responses seem harmful. Petersson et al. [78] and Jung et al. [49] describe a suspicion among users of AI chatbots for mental health support that data could be used for purposes that violate their privacy, and users describe intentionally anonymizing aspects of their experience out of fear that their privacy might be compromised. Hill [43] has described how publishing news reports about experiences of perceived sentience among chatbot users has led to numerous people reaching out to tech journalists to describe their own experiences, and describing experiences feeling unheard by AI companies. Adler [2] describes how, after an individual felt as if he had been misled by ChatGPT into delusions of saving the world, he attempted to “file a report to OpenAI so that they [could] fix ChatGPT’s behavior for other users.” However, this process was not fruitful, including false information from ChatGPT’s interface making it seem as if a report had been filed on the individual’s behalf, and after contacting OpenAI’s support team, boilerplate responses that insisted that he modify his personalization settings in the ChatGPT application or provided information about model hallucinations. Adler notes that this process left the individual with little clarity around whether “his reports had really been read or made their way to anyone with the power to change things.”
@@
A similarly opaque system globally is the process by which mental illness is diagnosed and treated, and treatment is subsequently paid for. In many systems, across international contexts, coverage of mental health services can often be tied to the diagnosis provided to a patient [57, 44], which may never be clearly discussed with the patient [68]. As a result, individuals may encounter situations where they receive services that they are unable to pay for, tied to a misdiagnosis or a documentation issue from their provider. Care authorization policy action (such as enabling the ability for an individual to appeal a decision internally, and ask for an independent external review [48, 35] if their services are not covered) ensure that individuals have a process of recourse when they feel like they have been unfairly treated, or what is sometimes called the “right to complain” [34].
@@
Like with care authorization, clear mechanisms for recourse after support experiences are distressing could be beneficial in the context of AI-mediated mental health support, particularly given the propensity for black-box model responses to be unexpected in harmful ways. A formal hub could be created (similar to that of The Human Line Project [103]) for individuals to report their experiences, with dedicated teams to investigate where model behavior went awry and report back to the individual why that behavior may have occurred. Following a reflective interpretability approach, including a clear and easily usable recourse mechanism would enable users to actively make sense of unexpected outputs after the fact, provide clarity on the boundaries that may have been crossed during their interaction, and give users a pathway to derive meaning from their distressing interaction, through working to make future responses more positive for other users via their report.
5 Tensions and Potential Risks
In the previous sections, we describe how reflective interpretability may allow a user to feel a greater sense of agency as they engage with AI-mediated mental health support, including providing more information to users around how responses are created and boundaries with regards to the kind of support that can be provided. However, multiple studies [96, 92, 49] have described how a major incentive for engaging with AI-mediated mental health support was the frictionless ease with which users felt able to find support from general purpose chatbots. With this in mind, it is important to consider the potential tensions and risks associated with the addition of additional design features (which could be experienced as more friction, and disincentivize engaging with AI-mediated mental health support).
For instance, as part of the intervention titration process, directly exposing intermediate CoT reasoning traces to users might be overwhelming and make the user feel pathologized in a way that they find hurtful or offensive. In clinical settings, clinicians are encouraged to reflect privately on how their experiences might influence their treatment planning for a client (a concept commonly called countertransference [107]). Intermediate chatbot reasoning traces that openly share perspectives on the user (such as “my next step is to persuade the user to shift their beliefs”) might make clients feel manipulated. Instead, the aim of a reflective interpretability approach is for responses to be part of “collaborative empiricism” [25], providing users enough (well-paced and well-timed) insight to preserve agency without granular information that overwhelms, deceives, or coerces them. It is thus important that reasoning traces are provided to users via interface design that highlights when model outputs are tailored to specific user customizations, versus when outputs follow a guiding therapeutic value or safety principle (especially when customizations are detected to be harmful), encouraging users to reflect on what next steps might most support their growth.